OpenStudy (anonymous):

Find the inverse of the function. f(x) = the cube root of quantity x divided by seven. - 9

OpenStudy (anonymous):

Can you write the equation for f (x) ?

OpenStudy (anonymous):

what do you mean ?

OpenStudy (anonymous):

Is the equation given as above as a description or are you just describing the equation?

OpenStudy (anonymous):

f-1(x) = 21(x + 9) f-1(x) = [7(x + 9)]3 f-1(x) = 7(x3 + 9) f-1(x) = 7(x + 9)3 these are my answer choices

OpenStudy (anonymous):

f(x) = the cube root of quantity x divided by seven. - 9 its the equation

OpenStudy (anonymous):

"the cube root of quantity x divided by seven. - 9" Can you write this as an equation? First step to finding inverse.

OpenStudy (the_fizicx99):

I hate word problems, >.> 

OpenStudy (anonymous):